OpenStudy (johan14th):

How to solve y'=y+x(y')+2x I'm having trouble moving the y'. Thanks :)

OpenStudy (johan14th):

I have tried subtracting xy' but that cancels out the y'

OpenStudy (johan14th):

My teacher some how got 2(dy^2/dx)=(y+2x)/2

OpenStudy (johan14th):

Sorry my teacher divided by 2 on bith sides

OpenStudy (whpalmer4):

\[y' = y + x y' + 2x\]\[y' - x y' = y + 2x\]\[y'(1-x) = y+2x\]is that what you needed?

OpenStudy (johan14th):

Okay thank you i was just thinking about it wrongly. This is what I needed thank you